OpenStudy (kdhdjhio):

I'll fan and medal Write the quadratic function in vertex form. y = x2 - 16x + 66

OpenStudy (kdhdjhio):

@TheSmartOne @agent0smith

jimthompson5910 (jim_thompson5910):

\[\Large y = x^2 - 16x + 66\] is the same as \[\Large y = 1x^2 +(-16x) + 66\]

jimthompson5910 (jim_thompson5910):

\[\Large {\color{red}{1}}x^2+({\color{green}{-16}}x)+{\color{blue}{66}}\] is in the form \[\Large {\color{red}{a}}x^2+{\color{green}{b}}x+{\color{blue}{c}}\] where \[\Large \color{red}{a = 1}\] \[\Large \color{green}{b = -16}\] \[\Large \color{blue}{c = 66}\]

jimthompson5910 (jim_thompson5910):

do you follow so far?

OpenStudy (kdhdjhio):

Yes @jim_thompson5910

jimthompson5910 (jim_thompson5910):

What we do from here is plug the values of 'a' and 'b' into \[\Large h = \frac{-b}{2a}\] So, \[\Large h = \frac{-(-16)}{2*1}\] \[\Large h = ???\] Tell me what you get

OpenStudy (kdhdjhio):

8

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

this means the x coordinate of the vertex is x = 8 plug x = 8 back into the original equation y = x^2 - 16x + 66 to find the corresponding value of y. Tell me what you get.

OpenStudy (kdhdjhio):

2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

when x = 8, the corresponding paired value of y is y = 2 the vertex is therefore (h,k) = (8,2) which means h = 8, k = 2 now we know the following a = 1 h = 8 k = 2 which we then plug into the vertex form template \[\Large y = a(x-h)^2 + k\]

OpenStudy (kdhdjhio):

I got x^2-16x +66

jimthompson5910 (jim_thompson5910):

So, \[\Large y = a(x-h)^2 + k\] turns into \[\Large y = 1(x-8)^2 + 2\] which simplifies to \[\Large y = (x-8)^2 + 2\]

jimthompson5910 (jim_thompson5910):

Don't expand out and combine like terms. Otherwise, you'll just get back to the original equation again.

OpenStudy (kdhdjhio):

Oh..Okay

jimthompson5910 (jim_thompson5910):

The final answer is \[\Large y = (x-8)^2 + 2\]

OpenStudy (kdhdjhio):

Thank you for helping, can you help with another?

jimthompson5910 (jim_thompson5910):

Sure I can help. Please close this current question and start a new thread. It's best to have one question per thread.